OpenStudy (anonymous):

Factor by first looking for a greatest common factor. 8z^2+24z+18

OpenStudy (anonymous):

First factor out 2 to get \[2(4x^{2}+12x+9)\]=\[2(4x^2+6x+6x+9)=2(2x(2x+3)+3(2x+3))=2(2x+3)^2\]